Find the sum of:
(a) (- 7) + (-9) + 4 + 16
(b) (37) + (-2) + (-65) + (-18)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) (- 7) + (-9) + 4 + 16 = (-7) + (-9) + 4 + (+ 7) + (+ 9) = (-7) + (+7) + (-9) + (+9) + 4 = 0 + 0 + 4 = 4 [∵ (-a) + (a) = 0] (b) (37) + (-2) + (-65) + (-8) = (+37) + (-75) = (+37) + (-37) + (-38) = 0 + (-38) = (-38) [∵ (-a) + (+a) = 0]